Many IRA fund managers argue that investors should invest at the beginning of the year rather than at the end. What is the difference to an investor who invests $2,000 per year at 11 percent over a 30 year period? Excel hint: Type = 0; ordinary annuity Type = 1: annuity due
Answer #1

There is a difference between investing fund at end of year and at beginning of year.Same is shown as under :

When an investor invest at end of year

FV(annuity) = Annuity[(1+r)^n-1 / r]

Annuity = $2000 , r = 11% , n =30 years

FV(annuity) = 2000[(1+11%)^30 - 1 /11%]

=2000[(1.11)^30-1 / 0.11]

=2000[(22.8923 - 1 /0.11]

=2000[21.8923/0.11]

=2000(199.0209)

=$398041.8

When an investor invest at beginning of year

FV(annuity) = Annuity[(1+r)^n-1 / r] (1+r)

=398041.8(1.11)

=441826.3$

Thus when one invest at beginning of year one get extra 441826.3-398041.8 = $43.784.59